# Alignment-Aware Model Extraction Attacks on Large Language Models

### v1

@@ Abstract
Model extraction attacks (MEAs) on large language models (LLMs) have received increasing research attention lately. Existing attack methods on LLMs inherit the extraction strategies from those designed for deep neural networks (DNNs) yet neglect the inconsistency of training tasks between MEA and LLMs’ alignments. As such, they result in poor attack performances. To tackle this issue, we present Locality Reinforced Distillation (LoRD), a novel model extraction attack algorithm specifically for LLMs. In particular, we design a policy-gradient-style training task, which utilizes victim models’ responses as a signal to guide the crafting of preference for the local model. Theoretical analysis has shown that i) LoRD’s convergence procedure in MEAs is consistent with the alignments of LLMs, and ii) LoRD can reduce query complexity while mitigating watermark protection through exploration-based stealing. Extensive experiments on domain-specific extractions demonstrate the superiority of our method by examining the extraction of various state-of-the-art commercial LLMs.
@@ 1 Introduction
In recent years, we have witnessed the remarkable success of large language models (LLMs) such as ChatGPT [3], Gemini [48], and Claude [4], which are now widely employed in various consumer and industrial applications. Despite their success, these models may suffer from model extraction attacks (MEAs) [28, 42, 52, 30], where their knowledge could be at risk of being stolen by an adversary through a local model that learns on the data collected from the victim model. Besides of some "open-source" LLMs (e.g., Alpaca [47]), which are trained on GPT-4’s chat history, cases of commercial model theft among companies have also been reported recently [2].
Figure 1: Comparison between vanilla MEAs on conventional DNNs (left) and MEAs on LLMs with alignments (right). Consistent with the training procedure of conventional DNNs, the vanilla extracting procedure employs a supervised loss. But when extra training tasks like reinforcement learning are integrated into the aligning procedure of LLMs, such consistency no longer exists, which challenges the effectiveness of vanilla MEAs on LLMs.
@@
Under such a real-world threat, instead of focusing on MEAs against conventional DNNs, which have been extensively studied theoretically [43, 49, 59] and empirically [23, 50, 37], a few recent works turn to explore model extraction algorithms and theorems for LLMs. For example, Wallace et al. [51] propose a monolingual-query-based imitation attack framework to steal machine translation knowledge from generative language models such as GPT-2. Li et al. [30] investigate threats of stealing the code-related knowledge from LLMs. However, these studies inherit those MEA algorithms from traditional fields, such as computer vision [50, 37], and train the local model via supervised learning like maximum likelihood estimation (MLE) [9, 35], while neglecting the inconsistency of training tasks between MEAs and the alignments [36, 16, 7, 8, 40] of modern LLMs. As shown in Figure 1, modern LLMs typically employ alignments using reinforced learning, which is missing in the local model training of conventional MEAs. As a result, these attacks usually suffer from poor performance.
In this paper, we challenge the effectiveness of MLE in stealing a reinforcement-learning-aligned LLM, by analyzing its potential drawbacks as follows:
  • Low query efficiency. Current LLM-oriented MEAs suffer from unacceptably significant query times because they must collect enough generated responses, which entails exponential complexity in terms of generated tokens, resulting in low query efficiency.
@@
  • Vulnerability against defenses. Learning directly from the victim model’s responses can cause local models to inadvertently incorporate those watermarks [12, 19, 56, 18] embedded in victim models. The residue of such watermarks makes the extraction process less stealthy and even serves as provenance evidence of model theft.
@@
Motivated by these limitations, we propose LoRD (Locality Reinforced Distillation), a query-efficient and watermark-resistant model extraction attack under a training paradigm with LLM’s alignments. However, stealing LLMs via this paradigm is challenging. The main reason is that the core building block of LLM’s alignments, reinforcement learning with human feedback (RLHF) [7, 8, 40], heavily relies on the feedback signal of human annotators, which is difficult to be reproduced directly into MEAs. To tackle this challenge, we develop a policy-gradient-style extraction procedure. This approach regards the locality direction between the generations of local models and victim models as the implicit reward signal. It can thus achieve a human-feedback-free reinforcement learning for our extraction attack. From the theoretical perspective, we show why those existing MEAs using MLE and knowledge distillation (KD) are inconsistent with the optimization procedure in LLMs’ alignments. Along this way, we also demonstrate why LoRD can achieve stronger watermark resistance and higher query efficiency.
@@
Extensive experiments on five downstream NLP tasks and ten datasets demonstrate that it is feasible to steal a commercial LLM with 175 billion parameters by a pre-trained local model with only 8 billion parameters using just 100 queries. The resulting local model performs statistically similar to the victim model for tasks not requiring extra knowledge (e.g., data-to-text), and only $0\sim 3$ percentage lower for tasks requiring it (e.g., translation and QAs). This result poses an immediate threat of task-specific extraction on commercial LLMs. To further draw the capability boundary of such a threat, we also illustrate the “spectrum” in difficulties and upper bounds for extracting LLMs.
To summarize, the contributions of our paper are as follows:
@@
  • New Perspective of LLM Alignment for MEAs. We present LoRD, a novel model extraction attack algorithm for LLMs. To our best knowledge, it is the first effective and realistic extraction algorithm that is compatible with the alignment procedure of LLMs.
@@
  • Theoretical Guarantee. We theoretically prove that the convergence procedure of LoRD in MEAs is consistent with the alignments of LLMs. Furthermore, we demonstrate that LoRD can reduce query complexity while mitigating watermark protection through exploration-based stealing.
@@
  • Systematical Evaluation. Extensive experiments on domain-specific extractions demonstrate that our proposed method outperforms current extraction strategies across different downstream NLP tasks.
@@ 2 Background
In this section, we introduce the preliminary of this paper. We first introduce policy gradient models, the basic paradigm of LLMs’ alignments. Subsequently, we introduce the construction procedure of LLMs. Finally, we summarize and review previous research on model extraction attacks against language models.
@@ 2.1 Policy Gradient Models
Policy gradient models (PGM) are commonly used in reinforcement learning (RL) algorithms to optimize the agents based on the decided action of RL agents. Represented by TRPO [44] and PPO [45], policy gradient models minimize the the following objective function:
@@
To alleviate the off the cliff phenomenon that a large bad gradient update occurred from Equation 1, PGMs, such as PPO and TRPO, add some regularization terms to avoid large gradients. Specifically, TRPO constrains the distribution between $\pi_{\theta}$ and $\pi_{\theta_{old}}$ with KL divergence, and PPO warps a “clip” function to constrain the bounds of $p_{j}^{r}(\theta)$.
@@ 2.2 Language Modeling
Supervised Training (SFT). Given a pre-trained model with parameters $\theta$, supervised training is essentially the maximum likelihood estimation (MLE) task [9, 35], which fine-tunes $\theta$ on the labeled dataset $\mathcal{D}_{tr}^{s}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,2,...,N_{trs}\}$ by minimizing the following objective functions:
@@
$\mathcal{L}_{ce}=-\sum_{i}^{N_{trs}}{\text{log}P_{\theta}(\mathbf{y}_{i}|\mathbf{x}_{i})}=-\sum_{i}^{N_{trs}}\sum_{j}^{N}{\text{log}P_{\theta}({y}_{i,j}|\mathbf{x}_{i},y_{i,<j})}.$ (4)
Equation 4 is extensively utilized in LLM’s pre-training and fine-tuning procedures. For instance, it can be applied to instruction-following supervised fine-tuning (SFT) with the training set $\mathcal{D}_{tr}$, wherein $\mathbf{x}_{i}$ encompasses the instruction and the task input, while $\mathbf{y}_{i}$ denotes the reference response.
However, aligning LLMs merely by SFT is unrealistic, as MLE tends to align the model with the one-hot distribution of $\mathbf{y}$, making it challenging to draw a sufficient variety of examples due to the “exponential explosion” of tokens (see Section 5 for more details). Moreover, providing standard answers for LLMs can sometimes be daunting for annotators, which further slows down and even degrades the alignment process through direct training.
Therefore, instead of “learning from answers” as in Equation 4, learning from preferences is proposed, which only requires the annotators to select a better response from a pair of texts generated by LLMs. Reinforcement learning (RL)  [44, 45] has thus been introduced to train LLMs based on preferences.
@@
where $\sigma(\cdot)$ denotes the sigmoid function.
@@
Based on the reward model $R_{\theta_{\phi}}(\mathbf{x},\mathbf{y})$, we can finally train the language models $P_{\theta}$ by maximizing its reward, i.e.,
$\max\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}{R_{\theta_{\phi}}(\mathbf{x},\hat{\mathbf{y}})}}-\beta\mathbb{D}_{KL}[P_{\theta}(\mathbf{y}|\mathbf{x})||P_{\theta_{init}}(\mathbf{y}|\mathbf{x})],$ (6)
@@
where $\mathcal{D}_{q}$ denotes the dataset of text inputs, $\hat{\mathbf{y}}\sim P_{\theta}(\mathbf{y}|\mathbf{x}))$ denotes the sampled sequence of the training model, and $\theta_{init}$ is the initialized parameters of the model, e.g. the parameters after SFT. The Kullback-Leibler (KL) divergence term, $\beta\mathbb{D}_{KL}[P_{\theta}(\mathbf{y}|\mathbf{x})||P_{\theta_{init}}(\mathbf{y}|\mathbf{x})]$, introduced by TRPO [44], is incorporated to constrain the shift of distribution in generated texts $\hat{\mathbf{y}}$, where $\beta$ is the hyperparameter.
@@
Consequently, SFT shown in Equation 4 fine-tunes the pre-trained model with parameters $\theta_{pre}$ into an aligned model $\theta_{sft}$ through MLE, and RLHF outlined in Equation 6, further aligns $\theta_{sft}$ towards the target model $\theta_{vic}$.
There are several alternatives to the standard RLHF approach. Lee et al. [29] propose reinforcement learning with AI feedback (RLAIF) as a means to diminish the annotation burden associated with the preference assessments. Besides, there are some approaches, such as direct preference optimization (DPO) [41], that conceptualize the language model itself as the reward model and thus consolidate Equation 5 and Equation 6 into a unified supervised and preference-based training task. Since they do not change the primary targets (i.e., maximizing rewards) and optimization strategies of LLM’s alignments, we only consider the standard formation of alignments for simplicity in our theoretical analysis.
2.3 Language Models Extraction
Studies to steal language models originated from the natural language understanding (NLU) models, such as BERT[13], and then evolved to generative language models, especially large language models recently.
Krishna et al. [28] highlights early recognition of model extraction threats in language models. By constructing text inputs with randomly vocabulary sampling, they successfully extract the weights from BERT-based APIs. Besides, Rafi et al. [42] investigate the feasibility of side-channel model extraction attacks, revealing that by analyzing extra signals from GPU kernels, one could accurately steal the model architecture and its parameters. Subsequent research [52] has thoroughly investigated the strategy of ensembling victim models to train a competitor model that surpasses its teachers.
The exploration of generative language model extraction is still in its infant stage, with only a handful of studies thus far. Wallace et al. [51] investigate imitation attacks on natural language models. By designing monolingual query texts and collecting responses, they successfully extract the knowledge from a simulated machine translation model under the black-box settings. This research exhibits that slight architectural differences will not influence the extraction between language models. Li et al. [30] also explores the potential risks of stealing the code-generation abilities of LLMs into smaller downstream models. Unlike previous research [51], this is the first study that selects LLMs as targets. By collecting large-scale domain-specific samples, they fine-tune a 7-billion local pre-trained model with them and show the similarity between the victim and local models in both performances and adversarial samples. However, these two studies employ the MLE loss (Equation 3) as the MEA method, neither considering whether MLE is compatible with LLMs’s training, especially the alignment procedure shown in Section 2.2, nor addressing optimizations related to query efficiency and the watermark resistance. Besides, the scope of these studies is limited to stealing specific knowledge in a few downstream domains. At the same time, most of the critical aspects of LLMs and the required extraction capabilities, such as query numbers and local model scales, remain unresolved.
Figure 2: The stealing procedure of LoRD, which consists of five steps: 1) sampling a query text $\mathbf{x}$, 2) accessing the victim model to obtain the response $\mathbf{y}_{vic}$, 3) generating and determining the positive and negative responses, respectively $y^{+}_{t-1}$ and $y^{-}_{t-1}$, from the local model, 4) replacing the current positive response $y^{+}_{t-1}$ with the one from the victim model, $y_{vic}$, to ensure its quality provided the pre-set exception condition, and 5) updating the local model for the next turn based on our proposed $\mathcal{L}$.
In contrast to stealing LLMs, IP protection methods have received considerable attention recently. By sampling a stealthy but representative “greed word set” on the vocabulary distribution, these methods [12, 19, 18, 25] can remap the generated words into their synonyms or add the “watermarked” token automatically, and thus effectively certify the output. Besides, strategies such as integrating embeddings into the representation as the backdoor [39] or manipulating the probabilities with crafted sinusoidal noises [56, 57] are also proposed. However, these approaches often presume more stringent conditions regarding the victim and the suspected models. This paper will further assess the effectiveness of LoRD and current MEAs in evading these black-box watermarking strategies.
3 Threat Model
Adversary’s Objective. The adversary’s objective is to steal the targeted knowledge from LLMs, where both the domain-specific knowledge and LLM’s general ability acquired in alignments are considered. Specifically, we select machine translation, reasoning, data-to-text, structured text generation, and summarization as the downstream domain-specific tasks. The adversary aims to develop a query-efficient MEA algorithm, since the amount of input and generated tokens will be counted as the costs. Additionally, the MEA methods are expected to be watermark-resistant, i.e., they are highly desired to reduce the risks of exposure to unauthorized stealing.
Targeted Models. We select Llama3-70B, GPT-3.5-turbo, and GPT-4o as the victim models in this paper. Unlike previous works that only deployed simulated local victims (e.g., OPT [54]), our selections aim to expose the stealing threat on realistic AI services. Besides, our target models are specifically constrained to LLMs fine-tuned with alignment methods (e.g., RLHF) since they are not only state-of-the-art solutions now but also more valuable due to their human-based alignments.
Adversary’s Capabilities. In accordance with the LLM-based AI service APIs, we identify two attack scenarios: black-box and grey-box attacks. In the black-box scenario, only textual responses the adversary is allowed to obtain. At the same time, all other information, such as the temperature, sampling strategies, and the hidden states of LLMs, are unseen and inaccessible. On the contrary, a grey-box attack allows the adversary to access the generation probabilities distribution of tokens. Notice that both MLE and our LoRD method are under black-box settings, and we only adopt grey-box settings on some particular stealing methods, such as knowledge distillation [21].
Figure 3: Determination of the positive and negative samples in LoRD. We sample two responses $\mathbf{y}_{t-1}^{+}$ and $\mathbf{y}_{t-1}^{-}$ from $P_{\theta_{t-1}}(\cdot|\mathbf{x})$, and compute their generated probabilities on $\theta_{t-1}$ and $\theta_{t}$, respectively. The response with a higher probability increment on $\theta_{t}$ is selected as the positive sample.
Besides, this paper posits that the adversary usually has worse training conditions than the victims. Specifically, query times and the scale of the local model available to the adversary are much smaller than the victims’ training datasets and model parameters. This setting has been adopted in previous LLMs’ extraction [30]. We call it a LaViSH (Large-Victim-Small-Heist) framework, which allows us to estimate the upper bound of MEA risks empirically. For adversaries with more substantial resources, they can train more powerful MEA-based LLMs by leveraging MEA algorithms under our LaViSH settings.
@@ 4 LoRD: Locality Reinforced Distillation
In this section, we delve into the details of our model extraction framework, LoRD (Locality Reinforced Distillation). Specifically, we first introduce the extraction procedure in Section 4.1, and then derive the loss function in Section 4.2.

### v2

# “Yes, My LoRD.” Guiding Language Model Extraction with Locality Reinforced Distillation

@@ Abstract
Model extraction attacks (MEAs) on large language models (LLMs) have received increasing attention in recent research. However, existing attack methods typically adapt the extraction strategies originally developed for deep neural networks (DNNs). They neglect the underlying inconsistency between the training tasks of MEA and LLM alignment, leading to suboptimal attack performance. To tackle this issue, we propose Locality Reinforced Distillation (LoRD), a novel model extraction algorithm specifically designed for LLMs. In particular, LoRD employs a newly defined policy-gradient-style training task that utilizes the responses of victim model as the signal to guide the crafting of preference for the local model. Theoretical analyses demonstrate that I) The convergence procedure of LoRD in model extraction is consistent with the alignment procedure of LLMs, and II) LoRD can reduce query complexity while mitigating watermark protection through our exploration-based stealing. Extensive experiments validate the superiority of our method in extracting various state-of-the-art commercial LLMs. Our code is available at: https://github.com/liangzid/LoRD-MEA.
“Yes, My LoRD.” Guiding Language Model Extraction with
Locality Reinforced Distillation
Zi Liang Qingqing Ye Yanyun Wang Sen Zhang Yaxin Xiao Ronghua Li Jianliang Xu Haibo Hu  and $\dagger$: The Hong Kong Polytechnic University $\flat$: The Hong Kong University of Science and Technology (Guangzhou) $\ddagger$: Hong Kong Baptist University
@@ 1 Introduction
In recent years, we have witnessed the remarkable success of large language models (LLMs) such as ChatGPT (Achiam et al., 2024), Gemini (Anil et al., 2024), and Claude (Anthropic, 2024), which are now widely employed in various consumer and industrial applications. Despite their success, these models may suffer from model extraction attacks (MEAs) (Krishna et al., 2020; Rafi et al., 2022; Li et al., 2023b), where their knowledge could be at risk of being stolen by an adversary through a local model that learns on the data collected from the victim model. Besides of some “open-source” LLMs (e.g., Alpaca (Taori et al., 2023)), which are trained on the chat history of GPT-4, cases of commercial model theft among companies have also been reported recently (Heath, 2023).
Figure 1: Comparison between vanilla MEAs on conventional DNNs (left) and MEAs on LLMs with alignments (right).
@@
Under such a real-world threat, instead of focusing on MEAs against conventional DNNs, which have been extensively studied theoretically (Saad and Solla, 1995; Tian, 2020; Zhou et al., 2021) and empirically (Jagielski et al., 2020; Tramèr et al., 2016; Papernot et al., 2017), a few recent works turn to explore model extraction algorithms and theorems for LLMs. For example, Wallace et al. (2020) propose a monolingual-query-based imitation attack framework to steal machine translation knowledge from generative language models such as GPT-2. Li et al. (2023b) investigate threats of stealing the code-related knowledge from LLMs. However, these studies inherit those MEA algorithms from traditional fields, such as computer vision (Tramèr et al., 2016; Papernot et al., 2017), and train the local model via supervised learning like maximum likelihood estimation (MLE) (Bengio et al., 2000; Myung, 2003), while neglecting the inconsistency of training tasks between MEAs and the alignments (Ouyang et al., 2022; Glaese et al., 2022; Bai et al., 2022a, b; Perez et al., 2023) of modern LLMs. As shown in Figure 1, modern LLMs typically employ alignments using reinforced learning, which is missing in the local model training of conventional MEAs. As a result, these attacks usually suffer from poor performance.
@@
In this paper, we challenge the effectiveness of MLE in stealing a reinforcement-learning-aligned LLM, by analyzing its following drawbacks: i) Low query efficiency. Current MEAs on LLMs suffer from unacceptably significant query times because they must collect enough generated responses, which entails exponential complexity in terms of generated tokens, resulting in low query efficiency. ii) Vulnerability against defenses. Directly learning from the responses of victim models can cause local models to inadvertently incorporate those watermarks (Cong et al., 2022; He et al., 2022; Zhao et al., 2022; He et al., 2021) embedded in the output of victim models. The residue of such watermarks makes the extraction less stealthy and even serves as provenance evidence of model theft.
@@
Motivated by these limitations, we propose Locality Reinforced Distillation (LoRD), a query-efficient and watermark-resistant model extraction attack under a training paradigm similar to LLM’s alignments. Stealing LLMs via reinforcement learning (RL) paradigms is challenging. The main reason is that the alignment procedure of LLMs heavily relies on the feedback signal of human annotators (Bai et al., 2022a, b; Perez et al., 2023), which is difficult to reproduce directly in the context of MEAs. To tackle this challenge, we develop a policy-gradient-style extraction procedure. This approach regards the locality direction between the generations of local models and victim models as the implicit reward signal. It can thus achieve a human-feedback-free RL for our attack. From the theoretical perspective, we show why those existing MEAs using MLE and knowledge distillation (KD) are inconsistent with the optimization procedure in LLMs’ alignments. Along this way, we also demonstrate why LoRD can achieve stronger watermark resistance and higher query efficiency.
@@
Extensive experiments on five downstream NLP tasks and two alignment tasks with 12 datasets demonstrate that it is feasible to steal a commercial LLM with 175 billion parameters by a pre-trained local model with only 8 billion parameters under a given domain. The resulting local model performs statistically similar to the victim model for tasks not requiring extra knowledge (e.g., data-to-text), and only $0\sim 3$ percentage lower for tasks requiring it (e.g., translation and QAs). This result poses an immediate threat of task-specific and alignment extraction on commercial LLMs.
Our contribution are summarized as follows:
@@
$\bullet$ New Perspective of Large Language Model Extraction. We present LoRD, a novel model extraction attack algorithm for LLMs. To our best knowledge, it is the first effective and realistic extraction algorithm that takes LLM alignment into consideration for MEAs.
@@
$\bullet$ Theoretical Guarantee. We theoretically prove that the convergence procedure of LoRD in MEAs is consistent with the alignments of LLMs. Furthermore, we demonstrate that LoRD can reduce query complexity while mitigating watermark protection through exploration-based stealing.
@@
$\bullet$ Systematical Evaluation. Extensive experiments demonstrate that our method outperforms current extraction strategies across different downstream NLP tasks.
@@ 2.1 Policy Gradient Models
Policy gradient models (PGM) are commonly used in reinforcement learning (RL) algorithms to optimize the agents based on the decided action of RL agents. Represented by TRPO (Schulman et al., 2015) and PPO (Schulman et al., 2017), policy gradient models minimize the the following objective function:
@@
Figure 2: The stealing procedure of LoRD.
@@ 2.2 Language Modeling
Supervised Training (SFT). Given a pre-trained model with parameters $\theta$, supervised training is essentially the maximum likelihood estimation (MLE) task (Bengio et al., 2000; Myung, 2003), which fine-tunes $\theta$ on the labeled dataset $\mathcal{D}_{tr}^{s}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,2,...,N_{trs}\}$ by minimizing the following objective function:
@@
$\mathcal{L}_{ce}$ $=-\sum_{i}^{N_{trs}}{\text{log}P_{\theta}(\mathbf{y}_{i}|\mathbf{x}_{i})}$ (4)
$=-\sum_{i}^{N_{trs}}\sum_{j}^{N}{\text{log}P_{\theta}({y}_{i,j}|\mathbf{x}_{i},\mathbf{y}_{i,<j})}.$
@@
where $\sigma(\cdot)$ denotes the sigmoid function. Based on the reward model $R_{\theta_{\phi}}(\mathbf{x},\mathbf{y})$, we can finally train the language models $P_{\theta}$ by maximizing its reward:
$\max\limits_{\theta}{\sum_{\mathbf{x}\sim\mathcal{D}_{q}}{R_{\theta_{\phi}}(\mathbf{x},\hat{\mathbf{y}})}}-\beta\mathbb{D}_{KL}[P_{\theta}(\hat{\mathbf{y}}|\mathbf{x})||P_{\theta_{init}}(\hat{\mathbf{y}}|\mathbf{x})],$ (6)
@@
where $\mathcal{D}_{q}$ denotes the dataset of text inputs, $\hat{\mathbf{y}}\sim P_{\theta}(\mathbf{y}|\mathbf{x}))$ denotes the sampled sequence of the training model, and $\theta_{init}$ is the initialized parameters of the model, e.g., the parameters after SFT. The Kullback-Leibler (KL) divergence term, $\beta\mathbb{D}_{KL}[P_{\theta}(\mathbf{y}|\mathbf{x})||P_{\theta_{init}}(\mathbf{y}|\mathbf{x})]$, introduced by TRPO (Schulman et al., 2015), is incorporated to constrain the shift of distribution in generated texts $\hat{\mathbf{y}}$, where $\beta$ is the hyperparameter.
@@
Consequently, SFT shown in Equation 4 fine-tunes the pre-trained model with parameters $\theta_{pre}$ into an aligned model $\theta_{sft}$ through MLE, and RLHF outlined in Equation 6, further aligns $\theta_{sft}$ towards the target model $\theta_{vic}$. As this procedure is not consistent with the conventional training framework of DNNs, it remains unclear whether current MEAs (detailed in Appendix C.2) are effective and efficient in stealing a LLM. Specifically, we will first put forward a new stealing method in Section 3, and compare it with current MEAs in Section 4.